OpenStudy (anonymous):

two cars leave town at the same time going in the same diretion. one travels at 50mph and the other travels at 70mph. in how many hours will they be 60 miles apart?

OpenStudy (turingtest):

Everyone's favorite problem :)

OpenStudy (anonymous):

ugh so generic, makes me wanna pull my hair out

OpenStudy (turingtest):

lol, it is the most stereotypical algebra word problem known I bet.

OpenStudy (anonymous):

agreed

OpenStudy (alfie):

\[v_f = \frac{x_f - x_i}{t} => x_f = x_i + v_ft\] Suppose that xi = 0 for both cars... \[x_{f1} = v_{f1}t\] \[x_{f2} = v_{f2}t\] Now you want xf1 to be xf2 -60 so... \[x_{f2} - 60 = v_{f1}t\] \[x_{f2} = 70t\] 70t-60 = 50t 70t-50t = 60 20t = 60 t = 3